OpenStudy (anonymous):

COULD SOMEONE PLEASE HELP ME! I DONT UNDERSTAND THIS QEUTION Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (lucaz):

standard form is: ax+by=c

OpenStudy (anonymous):

no its y - y1 = m(x - x1)

OpenStudy (anonymous):

i mean that how you solve it

OpenStudy (anonymous):

but im not sure how could you help me

OpenStudy (lucaz):

that is the point-slope form

OpenStudy (lucaz):

you have two given points, you just replace x, y and x1, y1 to find the slope of the line

OpenStudy (anonymous):

okay i know that part ... but i dont know how to slove for the answer i get confused

OpenStudy (lucaz):

y-y1/x-x1 = 5-6/1-4 = -1/-3 = 1/3 in point slope form: y-y1=m(x-x1) y-(6)=1/3(x-4) y=x/3-4/3+6 y=x/3+14/3 standard form is -x/3+y=14/3 or x/3-y=-14/3

OpenStudy (anonymous):

i see so when i get to the second step do i divide that numbers ?

OpenStudy (lucaz):

what step?

OpenStudy (anonymous):

o never mind i got it :)

OpenStudy (lucaz):

okay..

OpenStudy (anonymous):

thanks:)

OpenStudy (lucaz):

no problem